OpenStudy (anonymous):

find average rate of change as a reduced fraction the points are 0,3 and 7,5

OpenStudy (chaise):

The rate of change is the gradient and can be expressed as rise divide run. m = y-y/x-x m = 5-3/7-0 m = 2/7 I hope this helped

OpenStudy (anonymous):

it is distance walked in miles and number of hours spent walking

OpenStudy (chaise):

Well your answer will be 2/7 miles per hour pr approx 0.285 miles per hour :)